OpenStudy (anonymous):

Given the equation y - 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in standard form. -3/4x + y = 10 3x - 4y = -40 y = 3/4x + 12 my answer y = 3/4x + 10 @sleepyjess

OpenStudy (anonymous):

am i right?

OpenStudy (sleepyjess):

What is 8*.75?

OpenStudy (anonymous):

6

OpenStudy (sleepyjess):

Yep, now we have y - 4 = 3/4x + 6. How do we get 4 to the other side?

OpenStudy (anonymous):

oh so it is D

OpenStudy (sleepyjess):

Yep :)

OpenStudy (sleepyjess):

I like you, you are here to learn and not just to get answers.

OpenStudy (anonymous):

ok thanks i thought it was c lol

OpenStudy (anonymous):

yeah i hate pppl giving me answers

OpenStudy (anonymous):

doesnt help at all

OpenStudy (anonymous):

makes me feel bad too like i cheated n stuff

OpenStudy (sleepyjess):

that's good, I feel the same way :)

OpenStudy (anonymous):

i sumbited the review and it was c i guess you wren't supposed to distribute first or what? cause y-4 = 3/4(x+8) y-4+4 = 3/4x+8+4 y = 3/4x +12

OpenStudy (anonymous):

well they have the parenthesis there to seperate from x or something, but it's ok

OpenStudy (anonymous):

thnx so much

OpenStudy (anonymous):

:)

OpenStudy (sleepyjess):

hmmmm interesting.....

OpenStudy (anonymous):

wait it isn't that either?

OpenStudy (anonymous):

so weird..